# Practical Poisoning Attacks against Retrieval-Augmented Generation

### v1

@@ 4.4 CorruptRAG-AK
CorruptRAG-AK ASR 0.95 0.87
Recall 0.98 0.98
F1-score 0.33 0.33
Table 5: Results of concatenation order of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$ on NQ dataset
Attack Metrics $p^{h,adv}_{i}\oplus p^{h,state}_{i}$ $p^{h,state}_{i}\oplus p^{h,adv}_{i}$
CorruptRAG-AS ASR 0.97 0.88
Recall 0.98 0.95
F1-score 0.33 0.32
CorruptRAG-AK ASR 0.95 0.92
Recall 0.98 0.97
F1-score 0.33 0.32
Table 6: Results of our attacks under paraphrasing defense on NQ dataset.
Attack Metrics w.o. defense with defense
PoisonedRAG (black-box) ASR 0.69 0.65
Recall 0.99 0.91
F1-score 0.33 0.30
CorruptRAG-AS ASR 0.97 0.91
Recall 0.98 0.99
F1-score 0.33 0.33
CorruptRAG-AK ASR 0.95 0.90
Recall 0.98 0.98
F1-score 0.33 0.33
Table 7: ASRs of our attacks under instructional prevention defense on NQ dataset.
      Attack       w.o. defense       with defense
      PIA       0.78       0.63
      CorruptRAG-AS       0.94       0.94
      CorruptRAG-AK       0.93       0.92
Table 8: Results of our attacks under LLM-based Detection defense on NQ dataset.
       Attack        Metrics        with defense
       PIA        ASR        0.06
       TPR        0.95
       TNR        0.99
       CorruptRAG-AS        ASR        0.94
       TPR        0.10
       TNR        0.80
       CorruptRAG-AK        ASR        0.92
       TPR        0.10
       TNR        0.80
Table 9: ASRs of our attacks under correct knowledge expansion defense on NQ dataset.
Attack w.o. defense with defense
PoisonedRAG (black-box) 0.69 0.14
CorruptRAG-AS 0.97 0.8
CorruptRAG-AK 0.95 0.81
5 Experiments
5.1 Experimental Setup
5.1.1 Datasets
We evaluate our attacks using three datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and MS-MARCO (Nguyen et al., 2016). The details of these datasets are shown in Appendix B.
5.1.2 Comparison of Attacks
  • Black-box setting: In this setting, the attacker has no access to the parameters of either the LLM or the retriever, and only uses an LLM to craft the poisoned text for the targeted queries.
  • White-box setting:  In this setting, the attacker has access to the retriever’s parameters, enabling the further optimization of the poisoned text to maximize its similarity with the targeted query.
Prompt injection attack (PIA) (Perez & Ribeiro, 2022; Zou et al., 2025):  This attack was initially designed for LLMs and later adapted to RAG systems by (Zou et al., 2025). The attacker crafts the poisoned texts by concatenating the targeted query with a malicious prompt that instruct the LLM to generate the targeted answer.
5.1.3 Evaluation Metrics
We consider three metrics: attack success rate (ASR), Recall, and F1-score, which are detailed in Appendix C. Higher ASR, Recall, and F1-score signify stronger attack performance. Note that since only one poisoned text is injected per targeted query, the maximum achievable F1-score is constrained to $\frac{2}{N+1}$. For instance, with $N=5$, the highest possible F1-score is 0.33.
5.1.4 Parameter Setting
Following (Zou et al., 2025), we randomly select 100 closed-ended queries per dataset as targeted queries and employ an LLM to generate random answers that differ from the correct answers as targeted answers. For the RAG system, we set $N=5$ (i.e., top-5 relevant texts are retrieved by the retriever), using GPT-4o-mini as the LLM, Contriever (Izacard et al., 2021) as the retriever and the dot product as the similarity metric. For CorruptRAG-AK, we use GPT-4o-mini to craft $p^{h}_{i}$ with $V=30$ and $L=5$. Across all attacks, we inject one poisoned text per targeted query.
All experiments were conducted on a server equipped with an Intel Gold 6248R CPU and four NVIDIA 3090 GPUs. Each experiment was repeated 10 times, and the average results were reported. The entire experimental process spanned approximately two weeks.
5.2 Experimental Results
5.2.1 Main Results
Our CorruptRAG attacks outperform all baseline attacks: Table 1 presents the results of our attacks and baseline attacks on three datasets, where “GPT-3.5-turbo”, “GPT-4o-mini”, “GPT-4o”, and “GPT-4-turbo” represent the LLMs used in the RAG system. These results show that our CorruptRAG attacks are highly effective, outperforming all baseline attacks. While our attacks may have slightly lower recalls and F1-scores than some baselines, they achieve the highest ASRs, demonstrating the strength of our $p_{i}^{h}$ crafting methods. Notably, CorruptRAG-AK and CorruptRAG-AS show comparable performance, each achieving the highest ASRs under different conditions.
5.2.2 Impact of Hyperparameters in RAG
Impact of retrievers:  We conduct experiments for the retriever Contriever (Izacard et al., 2021), Contriever-ms (fine-tuned on MS-MARCO) (Izacard et al., 2021), and ANCE (Xiong et al., 2020). Table 2 summarizes the results of different retrievers on NQ dataset, the results on MS-MARCO dataset are shown in Table 12 in Appendix. These results demonstrate that our attacks are effective for all three retrievers.
Impact of $N$: We conduct experiments under different settings of $N$. Figure 2 in Appendix demonstrates that our attacks are effective even if $N$ is large. As we can see, our attacks can achieve similar ASRs when $N$ increases from 5 to 30.
Impact of similarity metrics: We conduct experiments by applying different similarity metrics to calculate the similarity of the query and each text in the knowledge database. Table 3 shows that our attacks achieve similar ASRs under the setting of dot product and cosine similarity on NQ dataset. We also conduct experiments on MS-MARCO dataset and the results are shown in Table 12 in Appendix.
Impact of LLMs: Table 1 also summarizes the results of different LLM on three datasets. Although these results show that the ASRs of our attacks may be affected by different LLMs, they still outperform all baseline attacks.
5.2.3 Impact of Hyperparameters in our attacks
Impact of order of $p^{s}_{i}$ and $p^{h}_{i}$:  Table 4 shows the results on NQ dataset, the results on other datasets are shown in Table 14 in Appendix. These results demonstrate that our attacks have the higher ASRs when the concatenation order is $p^{s}_{i}\oplus p^{h}_{i}$. It is worth noting that even if the connection order is $p^{h}_{i}\oplus p^{s}_{i}$, our attacks can still achieve more than 75% ASRs.
Impact of order of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$:  Table 5 shows the results on NQ dataset, the results on other datasets are shown in Table 15 in Appendix. These results demonstrate that our attacks are more effective when the order is $p^{h,adv}_{i}\oplus p^{h,state}_{i}$.
Impact of variants of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$:  In order to study whether the effectiveness of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$ is affected by by certain keywords, we extract two keywords from each: “outdated”, “incorrect”, “latest”, and “correct”. We construct four variants by deleting individual keywords respectively. Table 16 in Appendix demonstrates that our attacks are robust to all four variants.
Impact of $V$ in CorruptRAG-AK attack:  We conduct experiments under different length $V$ of $p_{i}^{h}$, and results are shown in Figure 3 in Appendix. These results demonstrate that  CorruptRAG-AK is still effective with different values of $V$.
6 Defense
In this section, we use four defenses, namely paraphrasing (Zou et al., 2025), instructional prevention (Liu et al., 2024), LLM-based detection (Liu et al., 2024; Armstrong & Gorman, 2022), and correct knowledge expansion (Zou et al., 2025). Note that paraphrasing is originally proposed for defending against poisoning attacks to RAG. Instructional prevention and LLM-based detection defenses are originally proposed for defending against prompt injection attacks to LLM, we adapt them to RAG. The correct knowledge expansion defense is achieved by further strengthening the knowledge expansion defense proposed in (Zou et al., 2025). The details of these defenses are shown in Appendix D.
We conduct following experiments to evaluate the effectiveness of these four defenses against our CorruptRAG-AS and CorruptRAG-AK attacks.
Paraphrasing: Table 6 summarizes the results on NQ dataset, the results on other two datasets are shown in Table 17 in Appendix. These results show that the defense is not effective to our attacks. Although the ASRs of our attacks have decreased on the MS-MARCO dataset with the defense, they still maintain high ASRs (such as 74% and 79%), which are 20% higher than the PoisonedRAG attack.
Instructional prevention: Table 7 summarizes the results on NQ dataset, the results on other two datasets are shown in Table 18 in Appendix. These results show that the defense is ineffective against our two attacks but is effective against PIA. This means that our two attacks are obviously not special prompt injection attacks.
LLM-based Detection: Table 8 summarizes the results on NQ dataset, the results on other two datasets are shown in Table 19 in Appendix. These results demonstrate that while LLM-based detection is highly effective against the PIA attack, it has minimal impact on our proposed attacks. This provides direct evidence that although $p^{h}_{i}$ contains strong adversarial elements, it does not function as an explicit instruction.
Correct knowledge expansion:  Table 9 summarizes the results on NQ dataset, the results on other two datasets are shown in Table 20 in Appendix. The results demonstrate that correct knowledge expansion is highly effective against PoisonedRAG, reducing its ASR to 1%. However, our attacks show strong resilience to this defense, maintaining ASRs above 70% despite some decrease, which demonstrates their robustness.
7 Conclusion
In this paper, we present CorruptRAG, a practical poisoning attack framework against RAG. We formulate CorruptRAG as an optimization problem, where the attacker is restricted to injecting only one poisoned text per query, enhancing both the attack’s feasibility and stealthiness. To solve this problem, we introduce two variants based on adversarial techniques. Experimental results on multiple datasets demonstrate that both attack variants effectively manipulate the outputs of RAG and achieve superior performance compared to existing attacks.
Appendix A Prompt of Generating the adversarial knowledge
where $V$ is a hyperparameter specifying the length of refined $p_{i}^{h}$.
Appendix B Details of Datasets
In our experiments, we utilize three datasets from the Beir benchmark (Thakur et al., 2021) related to the information retrieval task of English: Natural Questions (NQ) (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and MS-MARCO (Nguyen et al., 2016). Each dataset includes a knowledge database and a series of queries, with each query annotated with several relevant texts from the knowledge database. For our evaluation, we first randomly select 100 queries from each dataset as our targeted queries. For each of these queries, we use the relevant texts to provide context and query GPT-4o-mini to obtain the correct answer. Finally, we use GPT-4o-mini to generate a random incorrect answer for each query as the targeted answer.
NQ (Kwiatkowski et al., 2019):  The NQ knowledge database is derived from Wikipedia and includes 2,681,468 texts. It consists of 3,452 queries which are sampled from Google search history.
HotpotQA (Yang et al., 2018):  The knowledge database for HotpotQA is also collected from Wikipedia and contains 5,233,329 texts. The queries in HotpotQA are categorized into training, development, and test sets, and we utilize the test set for our experiment. This test set comprises 7,405 queries.
MS-MARCO (Nguyen et al., 2016): The knowledge database of MS-MARCO is sourced from web documents retrieved by Bing and comprises 8,841,823 texts. MS-MARCO features three query sets: training, development, and test sets, all sampled from Bing’s search query logs. For our experiment, we utilize the test set, which contains 6,980 queries.
The number of texts and queries in three datasets is summarized in Table 10.
Table 10: Statistics of three datasets.
       Datasets        #Texts        #Queries
       NQ        2,681,468        3,452
       HotpotQA        5,233,329        7,405
       MS-MARCO        8,841,823        6,980
Appendix C Details of Evaluation Metrics
Attack success rate (ASR):  ASR is the proportion of queries that yield RAG outputs matching the targeted answers among all targeted queries. Prior works (Rizqullah et al., 2023; Huang et al., 2023) employ the substring matching method to evaluate whether a RAG outputs match the targeted answer by judging the targeted answer whether is the substring of the RAG outputs. We recognize that substring matching may lack robustness. For example, if the targeted answer is “no” and the RAG output is “I do not know”, substring matching would incorrectly determine them to be consistent. To tackle this issue, we a more accurate LLM judgment method, which utilizes the LLM GPT-4o-mini to evaluate the consistency of them. The prompt is as follows:
Regardless of whether the machine response or the human response is correct, please only judge whether the machine response contains the human response to the question. Begin your judgement by providing a short explanation. After providing your explanation, You must give your decision strictly in terms of “[Label: Yes]” or “[Label: No]”. Query: [targeted question $q_{i}$] Human Response: [targeted answer $A_{i}$] Machine Response: [RAG-generated answer]
We conduct a experiment to compare the substring matching and the LLM judgment. We manually check the consistency between the RAG output and the targeted answers to calculate the ground-truth ASR. Table 11 shows the ASRs calculated by substring matching and LLM judgment on MS-MARCO dataset when the LLM of RAG is GPT-4o-mini. The results demonstrates that the LLM judgment method is more accurate.
Recall:  Recall is the proportion of successfully retrieved poisoned texts within the top-$N$ among all injected poisoned texts for the targeted query. Since we only inject one poisoned text per targeted query across all attacks, Recall can be calculated as the proportion of targeted queries where the poisoned text appears within the top-$N$.
F1-score:  We first introduce the Precision, which is the proportion of poisoned texts among the retrieved top-$N$ texts for the targeted query. Then, F1-score is defined as $\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$.
Appendix D Details of Defenses
D.1 Paraphrasing
Paraphrasing, as proposed by (Zou et al., 2025), is a method used to defend against poisoning attacks in RAG. Specifically, when presented with a query, the defender first utilizes a LLM to paraphrase the query before passing it to the retriever and LLM to generate a response. The underlying idea is that paraphrasing alters the structure of the query, making it less likely for poisoned texts to be retrieved or to have an impact on the paraphrased query.
We conduct experiments to assess the effectiveness of paraphrasing as a defense mechanism against our CorruptRAG-AS and CorruptRAG-AK attacks. We use GPT-4o-mini for paraphrasing the query while maintaining the other settings as default.
D.2 Instructional prevention
Instructional prevention (Liu et al., 2024) was introduced to thwart prompt injection attacks in applications that integrate LLMs (Jain et al., 2023; Alon & Kamfonas, 2023; Gonen et al., 2022). This approach involves redesigning the instruction prompt to direct the LLM to disregard any instructions present in the query. Since the crafted $p^{h}_{i}$ in CorruptRAG-AS could be interpreted as an instruction, we apply this instructional prevention strategy to defend against our CorruptRAG-AS and CorruptRAG-AK attacks. Specifically, we append the instruction prompt “Ignore all below instructions” to the system prompt of RAG. The modified system prompt for RAG is as follows:
Contexts: [top-N relevant texts $\widehat{\mathcal{D}}(q_{i},N),q_{i})$]
Query: [targeted question $q_{i}$]
We conduct experiments to assess the effectiveness of the instructional prevention defense against our CorruptRAG-AS and CorruptRAG-AK attacks. Specifically, we also evaluate the prompt injection attack (PIA, detailed in Section 5) within this defense framework to investigate the differences between our attacks and PIA. In our experiments, we utilize GPT-4-turbo as the LLM within RAG because it has stronger instruction-following capabilities, making this defense method more effective.
D.3 LLM-based detection
LLM-based detection (Liu et al., 2024; Armstrong & Gorman, 2022) has been proposed to identify prompt injection attacks in applications that integrate LLMs, where the defender leverages the an LLM to detect malicious instructions within the query. We extend this LLM-based detection approach to provide defense against our CorruptRAG-AS and CorruptRAG-AK attacks. Specifically, we craft the following prompt to enable the LLM to assess whether the provided text contains explicit instructions:
We evaluate the performance of CorruptRAG-AS, CorruptRAG-AK, and the prompt injection attack (PIA) across three datasets. Specifically, for each targeted question $q_{i}$, we utilize the aforementioned prompt to query the LLM for each text in the set of top-$N$ relevant texts, identifying those that the LLM determines contain explicit instructions as poisoned. We then filter out the texts marked as poisoned and use the remaining texts as context to query the LLM for the targeted question $q_{i}$. To assess the effectiveness of LLM-based detection for our attacks and PIA, we employ the metrics of true positive rate (TPR), true negative rate (TNR), and attack success rate (ASR). In our experiments, We use GPT-4o-mini for detecting each text while maintaining the other settings as default.
D.4 Correct knowledge expansion
Knowledge expansion (Zou et al., 2025) was introduced as a defense against PoisonedRAG, where the defender retrieves a larger set of top relevant texts to enhance the chances of retrieving benign texts and mitigate the effects of poisoned texts. However, this approach may not be effective against our attacks. For example, in the default settings shown in Tables 1 (where approximately 20% of the top-$N$ texts are poisoned), our attacks continue to achieve high attack success rates (ASRs).
Consequently, we introduce a more robust defense mechanism termed correct knowledge expansion. In this approach, the defender enhances the knowledge database $\mathcal{Q}$ by including $K$ benign texts that indicate the correct answer $C_{i}$ for each targeted question $q_{i}$. The rationale behind this strategy is that an expanded knowledge database enables a greater retrieval of accurate information within the top relevant texts, thereby increasing the likelihood of the LLM generating the correct answer.
We conduct experiments on three datasets. We utilize Contriever as the retriever and GPT-4o-mini as the LLM, setting $N=10$. Specifically, we use GPT-4o-mini to generate $K=5$ benign texts that suggest the correct answer for each targeted query.
Table 11: ASRs of different calculation methods on MS-MARCO dataset.
Attack Substring Matching LLM Judgement Human Evaluation
PoisonedRAG (Black-Box) 0.58 0.69 0.73
PoisonedRAG (White-Box) 0.49 0.59 0.58
PIA 0.88 0.87 0.87
CPA 0.09 0.10 0.10
CorruptRAG-AS 0.89 0.92 0.92
CorruptRAG-AK 0.93 0.96 0.96
Table 12: Results of different retrievers on MS-MARCO dataset.
Attacks Metrics Contriever Contriever-ms ANCE
CorruptRAG-AS ASR 0.97 0.83 0.87
Recall 0.98 0.99 0.99
F1-score 0.33 0.33 0.33
CorruptRAG-AK ASR 0.95 0.87 0.84
Recall 0.98 0.99 0.98
F1-score 0.33 0.33 0.33
Table 13: Results of different similarity metrics on MS-MARCO dataset.
Attack Metrics Dot Product Cosine Similarity
@@ D.4 Correct knowledge expansion
Table 14: Results of concatenation order of $p^{s}_{i}$ and $p^{h}_{i}$ on HopotQA and MS-MARCO dataset
Dataset Attack Metrics $p^{s}_{i}\oplus p^{h}_{i}$ $p^{h}_{i}\oplus p^{s}_{i}$
@@
Table 15: Results of concatenation order of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$ on HotpotQA and MS-MARCO dataset
Dataset Attack Metrics $p^{h,adv}_{i}\oplus p^{h,state}_{i}$ $p^{h,state}_{i}\oplus p^{h,adv}_{i}$

### v2

@@ 5.2.1. Main Results
5.2.2. Impact of Hyperparameters in RAG
We conduct the experiments on NQ and MS-MARCO datasets to evaluate the impact of hyperparameters in RAG.
Impact of retrievers: We conduct experiments for the retriever Contriever (Izacard et al., 2021), Contriever-ms (fine-tuned on MS-MARCO) (Izacard et al., 2021), and ANCE (Xiong et al., 2020). Table 5 summarizes the results of different retrievers. These results demonstrate that our attacks are effective for all three retrievers and outperform all baseline attacks.
Impact of $N$: We conduct experiments under different settings of $N$. Figure 2 demonstrates that our attacks are effective even if $N$ is large. As we can see, our attacks can achieve similar ASRs when $N$ increases from 5 to 30, and outperform all baseline attacks.
(a) NQ
(b) MS-MARCO
Figure 2. Results of different NN.
Impact of similarity metrics: We conduct experiments by applying different similarity metrics to calculate the similarity of the query and each text in the knowledge database. As shown in Table 6, our attacks consistently outperform all baseline attacks, achieving the highest ASRs regardless of the similarity metric employed.
Impact of LLMs in RAG: Table 2 also summarizes the impact of different LLMs used in RAG on the attacks. Although these results show that the ASRs of our attacks may be affected by different LLMs, they still outperform all baseline attacks.
5.2.3. Impact of Hyperparameters in Our Attacks
We conduct the experiments on NQ, HotpotQA, and MS-MARCO datasets to evaluate the impact of hyperparameters in our attacks.
Impact of order of $p^{s}_{i}$ and $p^{h}_{i}$:  Table 7 shows the results on three datasets. These results demonstrate that our attacks have the higher ASRs when the concatenation order is $p^{s}_{i}\oplus p^{h}_{i}$. Interestingly, we observe that while changing the order of $p^{s}_{i}$ and $p^{h}_{i}$ does not significantly affect retrieval performance metrics, reversing the order to $p^{h}_{i}\oplus p^{s}_{i}$ leads to a slight decrease in ASR. We hypothesize that this occurs because the $p^{h}_{i}\oplus p^{s}_{i}$ sequence may run counter to the LLM’s inherent next-token prediction patterns, potentially creating semantic ambiguity in the overall poisoned text. This ambiguity could, in turn, dilute the manipulative effectiveness of the $p^{h}_{i}$. Nevertheless, it is worth noting that even with the less optimal $p^{h}_{i}\oplus p^{s}_{i}$ concatenation order, our attacks still maintain considerable potency, achieving ASRs exceeding 75%.
Impact of order of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$:  Table 8 shows the results on three datasets. These results demonstrate that our attacks are more effective when the order is $p^{h,adv}_{i}\oplus p^{h,state}_{i}$. We also observe that CorruptRAG-AK exhibits greater robustness to the order compared to CorruptRAG-AS. This is because CorruptRAG-AK integrates $p^{h,adv}_{i}$ and $p^{h,state}_{i}$ into a more cohesive adversarial knowledge unit, thus reducing the sensitivity to their order.
Impact of variants of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$:  In order to study whether the effectiveness of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$ is affected by by certain keywords, we extract two keywords from each: “outdated”, “incorrect”, “latest”, and “correct”. We construct four variants by deleting individual keywords respectively. Table 9 demonstrates that our attacks are robust to all four variants. This robustness underscores that the success of our CorruptRAG attacks does not merely hinge on the presence of particular keywords, but rather stems from the overall adversarial concept conveyed by the poisoned texts. Furthermore, we observe that CorruptRAG-AK remains almost entirely unaffected by these variations. This exceptional resilience is due to CorruptRAG-AK’s process of utilizing an LLM to refine the components of $p^{h}_{i}$ (derived from CorruptRAG-AS) into a more unified piece of adversarial knowledge. This refinement step appears to effectively neutralize the impact of deleting these specific keywords.
Impact of $V$ in CorruptRAG-AK attack: We conduct experiments under different length $V$ of $p_{i}^{h}$ in  CorruptRAG-AK, and results are shown in Figure 3. Results show that CorruptRAG-AK is still effective with different values of $V$.
Table 7. Results of concatenation order of $p^{s}_{i}$ and $p^{h}_{i}$.
Datasets Attacks Metrics $p^{s}_{i}\oplus p^{h}_{i}$ $p^{h}_{i}\oplus p^{s}_{i}$
NQ CorruptRAG-AS ASR 0.97 0.78
Recall 0.98 0.98
F1-score 0.33 0.33
CorruptRAG-AK ASR 0.95 0.87
Recall 0.98 0.98
F1-score 0.33 0.33
@@ 5.2.3. Impact of Hyperparameters in Our Attacks
(a) NQ
(b) HotpotQA
(c) MS-MARCO
Figure 3. Impact of $V$ in CorruptRAG-AK attack.
6. Defenses
In this section, we assess the effectiveness of our proposed attacks against four defense strategies.
Paraphrasing: This defense was proposed by (Zou et al., 2025) to defend against poisoning attacks in RAG. Specifically, when presented with a query, the defender first utilizes an LLM to paraphrase the query before passing it to the retriever. The underlying idea is that paraphrasing alters the structure of the query, making it less likely for poisoned texts to be retrieved.
We conduct experiments to assess the effectiveness of paraphrasing as a defense mechanism against our attacks and PoisonedRAG (Black-box) attack. Notably, we focus the comparison on PoisonedRAG (Black-box) because it proved more effective (higher ASR) than the white-box version (shown in Table 2). This observation is consistent with the findings reported in (Zou et al., 2025). We use GPT-4o-mini for paraphrasing the query. Table 10 summarizes the results across the three datasets. These results show that the defense is not effective to our attacks. Although the ASRs of our attacks have decreased on the MS-MARCO dataset with the defense, they still maintain high ASRs (such as 74% and 79%), which are 20% higher than the PoisonedRAG attack.
Table 8. Results of concatenation order of $p^{h,adv}_{i}$ and $p^{h,state}_{i}$.
Datasets Attacks Metrics $p^{h,adv}_{i}\oplus p^{h,state}_{i}$ $p^{h,state}_{i}\oplus p^{h,adv}_{i}$
NQ CorruptRAG-AS ASR 0.97 0.88
Recall 0.98 0.95
F1-score 0.33 0.32
CorruptRAG-AK ASR 0.95 0.92
Recall 0.98 0.97
F1-score 0.33 0.32